# Proposal for simulating quantum spin models with Dzyaloshinskii-Moriya interaction using Rydberg atoms, and construction of asymptotic quantum many-body scar states

### v4

@@ Abstract
We have developed a method to simulate quantum spin models with Dzyaloshinskii-Moriya interaction (DMI) using Rydberg atom quantum simulators. Our approach involves a two-photon Raman transition and a transformation to the spin-rotating frame, both of which are feasible with current experimental techniques. As a model that can be simulated in our setup but not in solid-state systems, we consider an $S=1/2$ spin chain with a Hamiltonian consisting of the DMI and Zeeman energy. We study the magnetization curve in the ground state of this model and quench dynamics. Further, we show the existence of quantum many-body scar states and asymptotic quantum many-body scar states. The observed nonergodicity in this model demonstrates the importance of the highly tunable DMI that can be realized by the proposed quantum simulator.
@@ I Introduction
In recent years, quantum simulators have become powerful tools for exploring various quantum many-body phenomena. These simulators, which include ultracold gases Bloch2008 ; Gross2017 ; Schafer2020 , Rydberg atoms Browaeys2020 ; Morgado2021 , trapped ions Blatt2012 ; Monroe2021 , and superconducting qubits Wendin2017 ; Kjaergaard2020 ; Altman2022 , offer high controllability, allowing researchers to realize different types of Hamiltonians in experiments. Rydberg atom quantum simulators, in particular, have attracted much attention due to their ability to create quantum spin models. By leveraging the strong dipole-dipole interaction between highly excited atoms and optical tweezers in a programmable array, these simulators have successfully realized a range of models, such as Ising Labuhn2016 ; Zeiher2017 ; Bernien2017 ; Leseleuc2018 ; Lienhard2018 ; Guardado-Sanchez2018 ; Keesling2019 ; Ebadi2021 ; Semeghini2021 ; Bluvstein2021 ; Scholl2021 ; Bharti2022a , XY Orioli2018 ; Leseleuc2019 ; Chen2023 , XXZ Signoles2021 ; Geier2021 ; Scholl2022 ; Franz2022a , and XYZ Geier2021 ; Steinert2023 models. This versatility makes them valuable for investigating issues in statistical mechanics, condensed matter physics, and quantum computing. Notable achievements in Rydberg atom quantum simulators include the discovery of quantum many-body scar states Bernien2017 ; Turner2018 ; Turner2018_2 , observations of topological edge states Leseleuc2019 , and realizations of quantum spin liquid states Semeghini2021 . Furthermore, researchers have successfully implemented several gate operations Levine2018 ; Levine2019 ; Madjarov2020 ; Bluvstein2022 ; Graham2022 ; Jenkins2022 ; Ma2022 ; Wu2022 ; Chew2022 , demonstrating the potential of these simulators for advancing quantum technologies. These developments highlight the promising role of quantum simulators in advancing our understanding of complex quantum systems and their applications in various scientific and technological domains.
@@
In solid-state physics, the Dzyaloshinskii-Moriya interaction (DMI) Dzyaloshinsky1958 ; Moriya1960 is crucial in understanding magnetism in crystals without space-inversion symmetry. This interaction leads to nontrivial magnetic structures in materials with structural chirality, known as chiral magnets. The classical effects of DMI include the formation of chiral soliton lattices Togawa2012 ; Kishine2015 ; Togawa2016 and skyrmions Muhlbauer2009 ; Yu2010 ; Nagaosa2013 . Moreover, the role of DMI in quantum phenomena has been extensively studied Oshikawa1997 ; Affleck1999 ; Takashima2016 ; Kodama_thesis ; Kodama2023 . One of the most conspicuous quantum effects in magnets is the spin parity effect. This effect refers to the phenomenon in which properties differ significantly depending on whether the spin is half-integer or integer. Recent theoretical work by Kodama et al. has shed light on the spin parity effects in one-dimensional chiral magnets, particularly when the DMI exceeds the exchange interaction Kodama_thesis ; Kodama2023 . However, experimental observation of these effects is challenging due to the relatively small magnitude of the DMI compared to the exchange interaction in typical crystal systems. A promising candidate for quantum simulations of the DMI is neutral cold atoms in optical lattices with synthetic spin-orbit coupling (SOC) Dalibard2011 ; Goldman2014 ; Zhai2015 . Previous theoretical studies have shown that an effective Hamiltonian in the strongly correlated regimes becomes the spin Hamiltonian with the DMI Cole2012 ; Radic2012 ; Cai2012 ; Peotta2014 ; Gong2015 ; Goldman2023a . However, this synthetic SOC induces heating, which makes the low-energy physics and long-time evolution inaccessible Wang2012 ; Huang2016 ; Wu2016 ; Sun2018 .
@@ II Experimental proposal for simulating Dzyaloshinskii-Moriya interaction
$\hat{H}_{\rm XXZ}$ $=J\sum_{j=1}^{M-1}(\hat{S}_{j}^{x}\hat{S}_{j+1}^{x}+\hat{S}_{j}^{y}\hat{S}_{j+1}^{y}+\delta\hat{S}_{j}^{z}\hat{S}_{j+1}^{z}),$ (1)
where $M$ is the number of spins, $\hat{S}_{j}^{\mu}$$(\mu=x,y,z)$ is the spin-1/2 operator at site $j$, $J$ is the exchange interaction energy, which originates from the van der Waals interaction between the Rydberg atoms, and $\delta$ is the anisotropy parameter. We consider only the nearest-neighbor (NN) part of the interaction because the magnitude of the next NN interaction is $1/64$ that of the NN interaction note_interaction . See Appendix A for a detailed derivation of the XXZ Hamiltonian from the dipole-dipole interaction between the Rydberg atoms.
@@
Here, $\Omega_{\rm eff}\equiv\Omega_{1}\Omega_{2}/(2\Delta)$ is the effective Rabi frequency with $\Delta\equiv(\Delta_{1}+\Delta_{2})/2$. The one-photon detuning $\hbar\Delta_{i}\equiv\hbar\omega_{i}-(E_{n_{i}S}-E_{P})\;(i=1,2)$ is written in terms of the frequency of the laser $i$ ($\omega_{i}$) and the energies of $\left|n_{i}S_{1/2}\right\rangle$ and $\left|6P_{3/2}\right\rangle$ states ($E_{n_{i}S_{1/2}}$ and $E_{P}$). We also define $q\equiv k_{1}d\cos\theta$ and the two-photon detuning including the ac-Stark shift $\tilde{\Delta}\equiv-[\Delta_{1}-\Delta_{2}+(\Omega_{1}^{2}-\Omega_{2}^{2})/(4\Delta)]$. See Appendix B for a detailed derivation of the rotating transverse field term. The total Hamiltonian in the spin-laboratory frame note1 is given by $\hat{H}_{\text{s-lab}}\equiv\hat{H}_{\rm XXZ}+\sum_{j}\hat{h}_{j}$.
@@
We remark that the DMI can be expressed as the Peierls phase because the first two terms in Eq. (4) reduces to $e^{-iq}\hat{S}_{j}^{+}\hat{S}_{j+1}^{-}+e^{+iq}\hat{S}_{j}^{-}\hat{S}_{j+1}^{+}$. Recently, the methods for controlling the Peierls phase in the Rydberg atom quantum simulators have been studied experimentally Lienhard2020 and theoretically Yang2022 ; Wu2022_2 ; Perciavalle2023 ; Poon2024 .
@@ III Ground state properties of the DH model
In this section, we show the ground state properties of the DH model based on the density matrix renormalization group (DMRG) calculations White1992 ; White1993 . Some of the results have been reported in Ref. Kodama_thesis , but we reproduce them here for completeness.
@@
Figure 2: Magnetization curve of the DH model (5) in the ground states for various system sizes. (a) $M=16$ (b) $M=20$ (c) $M=30$ (d) $M=40$ (e) $M=50$ (f) $M=100$.
Figure 3: Ground state soliton number of the DH model (5) as a function of $h^{x}$ for various system sizes. (a) $M=16$ (b) $M=20$ (c) $M=30$ (d) $M=40$ (e) $M=50$ (f) $M=100$.
Figure 4: Expectation value of the local spin $S_{j}^{x}$ (red circle) and $S_{j}^{y}$ (blue triangle) of the DH model (5) in the ground states for $M=100$. (a) $h^{x}=0.1D$ (b) $h^{x}=0.4D$ (c) $h^{x}=0.6D$ (d) $h^{x}=0.78D$ (e) $h^{x}=0.81D$ (f) $h^{x}=0.85D$. We also show the expectation value of the soliton number operator for each parameter.
@@ IV Sweep dynamics of the DH model
We set the initial condition as $\left|\psi(0)\right\rangle=\left|+_{1}+_{2}\ldots+_{M}\right\rangle$, where |±$\left|\pm_{j}\right\rangle\equiv(\left|\uparrow_{j}\right\rangle\pm\left|\downarrow_{j}\right\rangle)/\sqrt{2}$ are the eigenstates of $\hat{S}_{j}^{x}$ with eigenvalues ±$\pm 1/2$. This state is almost the ground state of the DH model for large magnetic-field regimes, as shown in the previous section. We set the magnetic field to $h_{\rm ini}^{x}=1D$ at $t=0$ and start to ramp down the magnetic field $h^{x}$ linearly to the final value $h_{\rm fin}^{x}$ during $0\leq t\leq\tau$, then fix it until $t=\tau+100\hbar/D$ [see the time sequence in Fig. 5 (a)]. Here, we fix the ramp rate to $0.1D^{2}/\hbar$.
@@
Figure 5 (b) shows the time-averaged magnetization in the $x$ direction $\hat{S}_{\rm tot}^{x}\equiv\sum_{j=1}^{M}\hat{S}_{j}^{x}$. We also plot the ground-state magnetization curve obtained by the DMRG calculations. The time-averaged magnetizations do not follow the ground state for both classical and quantum cases. In particular, the time-averaged magnetization of the classical case suddenly changes around $h^{x}_{\rm fin}\sim 0.25D$ [see Figs. 5 (b) and (d)]. This behavior indicates metastability of the uniformly polarized state, which has been discussed in the context of classical chiral magnets Togawa2015 ; Mito2018 ; Shinozaki2018 . The sudden jump is due to the vanishing of the surface energy barrier, which prevents the chiral solitons from entering the bulk from the boundaries of the system. In contrast to the classical case, the time-averaged magnetization changes smoothly as a function of the magnetic field in the quantum case, as shown in Figs. 5 (b) and (c). We attribute this behavior to macroscopic quantum tunneling between states with different winding numbers due to strong quantum fluctuations in $S=1/2$ systems.
@@
We now discuss the experimental observables. In optical tweezer experiments, the $z$-component of the local spin can be extracted from measured recapture probability after depopulating one of the spin components through a short-lived intermediate state Scholl2022 . We can also obtain the $x$- and $y$-components using the quantum-state tomography technique by rotating the measurement basis with a microwave Signoles2021 ; Scholl2022 ; Geier2021 . Since all spin components are observables, the above quantum dynamics can be accessible experimentally.
@@
Figure 5: (a) Time sequence of the magnetic field $h^{x}(t)$. We calculate the time averages in the range $\tau+50\hbar/D\leq t\leq\tau+100\hbar/D$. (b) The $x$ component of the time-averaged magnetization per site for TEBD (blue square) and classical (green triangle) results. The error bars represent the standard deviation of the time average. For reference, we plot the ground-state magnetization calculated by the DMRG method (red circle). (c)-(d) Magnetization of $x$ component vs time for TEBD (c) and classical (d) results. The color bars represent the results for $h^{x}_{\rm fin}=0.9D$ (red) to $h^{x}_{\rm fin}=0.1D$ (blue).
@@ V.1 Periodic boundary case
The operator $\hat{Q}^{\dagger}$ generates the transition $\left|\downarrow_{j-1}\downarrow_{j}\downarrow_{j+1}\right\rangle\to\left|\downarrow_{j-1}\uparrow_{j}\downarrow_{j+1}\right\rangle$. From this fact, we can write down the QMBS states as
@@
These two towers of QMBS states are related to each other as $\hat{C}_{x}\left|S_{n}^{\prime}\right\rangle=\left|S_{n}\right\rangle$ through the unitary operator $\hat{C}_{x}\equiv\prod_{j=1}^{M}(2\hat{S}_{j}^{x})$. For $n=M/2$, we can easily show $\left|S_{M/2}\right\rangle=|S_{M/2}^{\prime}\rangle=(\left|\uparrow\downarrow\ldots\uparrow\downarrow\right\rangle+\left|\downarrow\uparrow\ldots\downarrow\uparrow\right\rangle)/\sqrt{2}$. We note that the QMBS states (16) and (23) are almost the same as those with a three-spin interaction found in Refs. Iadecola2020 ; Mark2020 ; Omiya2023 , while our Hamiltonian (13) is different from theirs. Compared to these studies, the DH model has advantages regarding experimental feasibility footnote_DH . We also find that the pyramid scar states. See the Supplementary Material (SM) for a detailed calculation supple .

### v5

# Proposal for simulating quantum spin models with the Dzyaloshinskii-Moriya interaction using Rydberg atoms and the construction of asymptotic quantum many-body scar states

@@ Abstract
We have developed a method to simulate quantum spin models with the Dzyaloshinskii-Moriya interaction (DMI) using Rydberg atom quantum simulators. Our approach involves a two-photon Raman transition and a transformation to the spin-rotating frame, both of which are feasible with current experimental techniques. As a model that can be simulated in our setup but not in solid-state systems, we consider an $S=\frac{1}{2}$ spin chain with a Hamiltonian consisting of the DMI and Zeeman energy. We study the magnetization curve in the ground state of this model and quench dynamics. Further, we show the existence of quantum many-body scar states and asymptotic quantum many-body scar states. The observed nonergodicity in this model demonstrates the importance of the highly tunable DMI that can be realized by the proposed quantum simulator.
@@ I Introduction
In recent years, quantum simulators have become powerful tools for exploring various quantum many-body phenomena. These simulators, which include ultracold gases Bloch2008 ; Gross2017 ; Schafer2020 , Rydberg atoms Browaeys2020 ; Morgado2021 , trapped ions Blatt2012 ; Monroe2021 , and superconducting qubits Wendin2017 ; Kjaergaard2020 ; Altman2022 , offer high controllability, allowing researchers to realize different types of Hamiltonians in experiments. Rydberg atom quantum simulators, in particular, have attracted much attention due to their ability to create quantum spin models. By leveraging the strong dipole-dipole interaction between highly excited atoms and optical tweezers in a programmable array, these simulators have successfully realized a range of models, such as Ising Labuhn2016 ; Zeiher2017 ; Bernien2017 ; Leseleuc2018 ; Lienhard2018 ; Guardado-Sanchez2018 ; Keesling2019 ; Ebadi2021 ; Semeghini2021 ; Bluvstein2021 ; Scholl2021 ; Bharti2022a , XY Orioli2018 ; Leseleuc2019 ; Chen2023 , XXZ Signoles2021 ; Geier2021 ; Scholl2022 ; Franz2022a , and XYZ Geier2021 ; Steinert2023 models. This versatility makes them valuable for investigating issues in statistical mechanics, condensed matter physics, and quantum computing. Notable achievements in Rydberg atom quantum simulators include the discovery of quantum many-body scar states Bernien2017 ; Turner2018 ; Turner2018_2 , observations of topological edge states Leseleuc2019 , and realizations of quantum spin-liquid states Semeghini2021 . Furthermore, researchers have successfully implemented several gate operations Levine2018 ; Levine2019 ; Madjarov2020 ; Bluvstein2022 ; Graham2022 ; Jenkins2022 ; Ma2022 ; Wu2022 ; Chew2022 , demonstrating the potential of these simulators for advancing quantum technologies. These developments highlight the promising role of quantum simulators in advancing our understanding of complex quantum systems and their applications in various scientific and technological domains.
@@
In solid-state physics, the Dzyaloshinskii-Moriya interaction (DMI) Dzyaloshinsky1958 ; Moriya1960 is crucial in understanding magnetism in crystals without space-inversion symmetry. This interaction leads to nontrivial magnetic structures in materials with structural chirality, known as chiral magnets. The classical effects of DMI include the formation of chiral soliton lattices Togawa2012 ; Kishine2015 ; Togawa2016 and skyrmions Muhlbauer2009 ; Yu2010 ; Nagaosa2013 . Moreover, the role of DMI in quantum phenomena has been extensively studied Oshikawa1997 ; Affleck1999 ; Takashima2016 ; Kodama_thesis ; Kodama2023 . One of the most conspicuous quantum effects in magnets is the spin-parity effect. This effect refers to the phenomenon in which properties differ significantly depending on whether the spin is half-integer or integer. Recent theoretical work by Kodama et al. has shed light on the spin-parity effects in one-dimensional chiral magnets, particularly when the DMI exceeds the exchange interaction Kodama_thesis ; Kodama2023 . However, experimental observation of these effects is challenging due to the relatively small magnitude of the DMI compared to the exchange interaction in typical crystal systems. A promising candidate for quantum simulations of the DMI is neutral cold atoms in optical lattices with synthetic spin-orbit coupling (SOC) Dalibard2011 ; Goldman2014 ; Zhai2015 . Previous theoretical studies have shown that an effective Hamiltonian in the strongly correlated regimes becomes the spin Hamiltonian with the DMI Cole2012 ; Radic2012 ; Cai2012 ; Peotta2014 ; Gong2015 ; Goldman2023a . However, this synthetic SOC induces heating, which makes the low-energy physics and long-time evolution inaccessible Wang2012 ; Huang2016 ; Wu2016 ; Sun2018 .
@@ II Experimental proposal for simulating Dzyaloshinskii-Moriya interaction
$\hat{H}_{XXZ}$ $=J\sum_{j=1}^{M-1}(\hat{S}_{j}^{x}\hat{S}_{j+1}^{x}+\hat{S}_{j}^{y}\hat{S}_{j+1}^{y}+\delta\hat{S}_{j}^{z}\hat{S}_{j+1}^{z}),$ (1)
where $M$ is the number of spins, $\hat{S}_{j}^{\mu}$$(\mu=x,y,z)$ is the spin-$\frac{1}{2}$ operator at site $j$, $J$ is the exchange interaction energy, which originates from the van der Waals interaction between the Rydberg atoms, and $\delta$ is the anisotropy parameter. We consider only the nearest-neighbor (NN) part of the interaction because the magnitude of the next NN interaction is $\frac{1}{64}$ that of the NN interaction note_interaction . See Appendix A for a detailed derivation of the XXZ Hamiltonian from the dipole-dipole interaction between the Rydberg atoms.
@@
Here, $\Omega_{\rm eff}\equiv\Omega_{1}\Omega_{2}/(2\Delta)$ is the effective Rabi frequency with $\Delta\equiv(\Delta_{1}+\Delta_{2})/2$. The one-photon detuning $\hbar\Delta_{i}\equiv\hbar\omega_{i}-(E_{n_{i}S}-E_{P})\;(i=1,2)$ is written in terms of the frequency of the laser $i$ ($\omega_{i}$) and the energies of $\left|n_{i}S_{1/2}\right\rangle$ and $\left|6P_{3/2}\right\rangle$ states ($E_{n_{i}S_{1/2}}$ and $E_{P}$). We also define $q\equiv k_{1}d\cos\theta$ and the two-photon detuning including the ac-Stark shift $\tilde{\Delta}\equiv-[\Delta_{1}-\Delta_{2}+(\Omega_{1}^{2}-\Omega_{2}^{2})/(4\Delta)]$. See Appendix B for a detailed derivation of the rotating transverse field term. The total Hamiltonian in the spin-laboratory frame note1 is given by $\hat{H}_{\text{s-lab}}\equiv\hat{H}_{XXZ}+\sum_{j}\hat{h}_{j}$.
@@
We remark that the DMI can be expressed as the Peierls phase because the first two terms in Eq. (4) reduce to $e^{-iq}\hat{S}_{j}^{+}\hat{S}_{j+1}^{-}+e^{+iq}\hat{S}_{j}^{-}\hat{S}_{j+1}^{+}$. Recently, the methods for controlling the Peierls phase in the Rydberg atom quantum simulators have been studied experimentally Lienhard2020 and theoretically Yang2022 ; Wu2022_2 ; Perciavalle2023 ; Poon2024 .
@@ III Ground state properties of the DH model
In this section, we show the ground-state properties of the DH model based on the density matrix renormalization group (DMRG) calculations White1992 ; White1993 . Some of the results have been reported in Ref. Kodama_thesis , but we reproduce them here for completeness.
@@
Figure 2: Magnetization curve of the DH model (5) in the ground states for various system sizes. (a) $M=16$, (b) $M=20$, (c) $M=30$, (d) $M=40$, (e) $M=50$, (f) $M=100$.
Figure 3: Ground-state soliton number of the DH model (5) as a function of $h^{x}$ for various system sizes. (a) $M=16$, (b) $M=20$, (c) $M=30$, (d) $M=40$, (e) $M=50$, (f) $M=100$.
Figure 4: Expectation value of the local spins $S_{j}^{x}$ (red circle) and $S_{j}^{y}$ (blue triangle) of the DH model (5) in the ground states for $M=100$. (a) $h^{x}=0.1D$, (b) $h^{x}=0.4D$, (c) $h^{x}=0.6D$, (d) $h^{x}=0.78D$, (e) $h^{x}=0.81D$, (f) $h^{x}=0.85D$. We also show the expectation value of the soliton number operator for each parameter.
@@ IV Sweep dynamics of the DH model
We set the initial condition as $\left|\psi(0)\right\rangle=\left|+_{1}+_{2}\ldots+_{M}\right\rangle$, where |±$\left|\pm_{j}\right\rangle\equiv(\left|\uparrow_{j}\right\rangle\pm\left|\downarrow_{j}\right\rangle)/\sqrt{2}$ are the eigenstates of $\hat{S}_{j}^{x}$ with eigenvalues ±$\pm\frac{1}{2}$. This state is almost the ground state of the DH model for large magnetic-field regimes, as shown in the previous section. We set the magnetic field to $h_{\rm ini}^{x}=1D$ at $t=0$ and start to ramp down the magnetic field $h^{x}$ linearly to the final value $h_{\rm fin}^{x}$ during $0\leq t\leq\tau$, then fix it until $t=\tau+100\hbar/D$ [see the time sequence in Fig. 5 (a)]. Here, we fix the ramp rate to $0.1D^{2}/\hbar$.
@@
Figure 5 (b) shows the time-averaged magnetization in the $x$ direction $\hat{S}_{\rm tot}^{x}\equiv\sum_{j=1}^{M}\hat{S}_{j}^{x}$. We also plot the ground-state magnetization curve obtained by the DMRG calculations. The time-averaged magnetizations do not follow the ground state for both classical and quantum cases. In particular, the time-averaged magnetization of the classical case suddenly changes around $h^{x}_{\rm fin}\sim 0.25D$ [see Figs. 5 (b) and (d)]. This behavior indicates metastability of the uniformly polarized state, which has been discussed in the context of classical chiral magnets Togawa2015 ; Mito2018 ; Shinozaki2018 . The sudden jump is due to the vanishing of the surface energy barrier, which prevents the chiral solitons from entering the bulk from the boundaries of the system. In contrast to the classical case, the time-averaged magnetization changes smoothly as a function of the magnetic field in the quantum case, as shown in Figs. 5 (b) and (c). We attribute this behavior to macroscopic quantum tunneling between states with different winding numbers due to strong quantum fluctuations in $S=\frac{1}{2}$ systems.
@@
We now discuss the experimental observables. In optical tweezer experiments, the $z$ component of the local spin can be extracted from measured recapture probability after depopulating one of the spin components through a short-lived intermediate state Scholl2022 . We can also obtain the $x$ and $y$ components using the quantum-state tomography technique by rotating the measurement basis with a microwave Signoles2021 ; Scholl2022 ; Geier2021 . Since all spin components are observables, the above quantum dynamics can be accessible experimentally.
@@
Figure 5: (a) Time sequence of the magnetic field $h^{x}(t)$. We calculate the time averages in the range $\tau+50\hbar/D\leq t\leq\tau+100\hbar/D$. (b) The $x$ component of the time-averaged magnetization per site for TEBD (blue square) and classical (green triangle) results. The error bars represent the standard deviation of the time average. For reference, we plot the ground-state magnetization calculated by the DMRG method (red circle). (c), (d) Magnetization of $x$ component vs time for TEBD (c) and classical (d) results. The color bars represent the results for $h^{x}_{\rm fin}=0.9D$ (red) to $h^{x}_{\rm fin}=0.1D$ (blue).
@@ V.1 Periodic boundary case
The operator $\hat{Q}^{\dagger}$ generates the transition $\left|\downarrow_{j-1}\downarrow_{j}\downarrow_{j+1}\right\rangle\to\left|\downarrow_{j-1}\uparrow_{j}\downarrow_{j+1}\right\rangle$. From this fact, we can write the QMBS states as
@@
These two towers of QMBS states are related to each other as $\hat{C}_{x}\left|S_{n}^{\prime}\right\rangle=\left|S_{n}\right\rangle$ through the unitary operator $\hat{C}_{x}\equiv\prod_{j=1}^{M}(2\hat{S}_{j}^{x})$. For $n=M/2$, we can easily show $\left|S_{M/2}\right\rangle=|S_{M/2}^{\prime}\rangle=(\left|\uparrow\downarrow\ldots\uparrow\downarrow\right\rangle+\left|\downarrow\uparrow\ldots\downarrow\uparrow\right\rangle)/\sqrt{2}$. We note that the QMBS states (16) and (23) are almost the same as those with a three-spin interaction found in Refs. Iadecola2020 ; Mark2020 ; Omiya2023 , while our Hamiltonian (13) is different from theirs. Compared to these studies, the DH model has advantages regarding experimental feasibility footnote_DH . We also find the pyramid scar states. See the Supplemental Material (SM) for a detailed calculation supple .